# Model Synthesis for Zero-Shot Model Attribution

### v2

@@ IV-D Comparison with Existing Methods (EQ3)
IV-E Model Lineage Analysis (EQ4)
@@
V Conclusions and Future Work
@@
Beyond the scope of this paper, there are several promising directions worth further explorations. Firstly, the generated images may undergo subsequent processing in practical scenarios, which can diminish their intrinsic spectrum patterns. It is important to develop methods for extracting robust patterns to prepare the fingerprint extractor for more practical applications. Secondly, aside from LoRA, there are many other types of model variants derived from a base model. Exploring more relationships between models through fingerprint extraction holds potential for researchers. Finally, while our method has shown effective generalization ability of training solely on synthetic models and tested on real models without fine-tuning, it does not outperform the performance of directly training on real models and test on these models. This indicates that the potential of model synthesizing has not been fully exploited. How to further increase the fidelity and diversity of model synthesis is also an important research direction in the future.

### v3

@@ IV-C Comparison with Existing Methods (EQ2)
TABLE X: Silhouette Score (SS) and Davies-Bouldin Index (DBI) comparison for fingerprint representations (higher values of Silhouette Score and lower values of Davies-Bouldin Index indicate better performance).
Method Classical Models Emerging Models
[…]
Marra et al. [5] -0.07 23.18 -0.21 34.17
Yu et al. [6] 0.36 1.21 -0.05 12.73
DNA-Det [8] 0.40 1.08 0.01 4.53
RepMix [9] 0.28 2.59 -0.10 8.48
Abady et al. [11] 0.38 1.18 -0.02 10.96
POSE [10] 0.44 0.90 -0.02 3.64
Ours (Only-Syn) 0.19 2.02 0.19 1.84
Ours (Fine-tune) 0.45 0.84 0.18 1.88
Evaluation of Fingerprint Representation. In Table X, we include additional evaluation metrics to assess the quality of fingerprint representation: 1)Silhouette Score (SS) [56], which measures how similar a point is to its own cluster compared to other clusters. Higher values indicate better separation between fingerprints of different models. 2) Davies-Bouldin Index (DBI) [57], which measures the ratio of intra-cluster distance to inter-cluster distance. Lower values indicate more compact and separated fingerprint representation. As shown in the table, existing methods such as DNA-Det and POSE achieve relatively high SS and low DBI on seen classical models. However, their generalization to unseen emerging models is limited, with SS decreasing and DBI increasing significantly. In contrast, our method, Ours (Only-Syn), demonstrates more consistent performance across both classical and emerging models. Notably, after fine-tuning on classical models, Ours (Fine-tune) achieves the highest SS and lowest DBI compared with existing methods, indicating superior fingerprint separation and robust generalization across diverse generative models.
IV-D Model Lineage Analysis (EQ3)
@@
IV-E Visualization Results (EQ4)
Real Models
Synthetic Models
Real Models
Synthetic Models
Figure 10: Spectrum of real-world models and their closest synthetic models within the 5760 synthetic models.
We provide spectrum visualizations of real models and their closest synthetic models in Figure 10. As shown, while the spectra of synthetic models do not exactly replicate those of real-world models, they exhibit notable similarities and diversity in frequency artifacts. This demonstrates the fidelity of our synthetic model generation process and its ability to approximate the spectral characteristics of real-world models effectively.
V Conclusions